OpenStudy (anonymous):

Change 4x – 3y = –15 to slope-intercept form. State slope and y-intercept.

OpenStudy (anonymous):

heeelpppp!

OpenStudy (anonymous):

4x+15=3y y=4x/5+5 y=mx+c m=slope=4/5 y intercept=5

OpenStudy (anonymous):

thanks dude

OpenStudy (anonymous):

giv me medal yar

OpenStudy (anonymous):

I gotchuuu!

OpenStudy (anonymous):

Are like medals good or something?

OpenStudy (anonymous):

just for fun yar thank u

OpenStudy (anonymous):

yeah man np